A nearsighted person has a far point of 50cm. What prescription(s) should the optometrist write in order to correct this person's vision?
Answer #1

here,

u = infinity

v = - 50 cm

using lens formula

1/f = 1/v - 1/u

1/f = - 1/50

power

P = 100/ f = - 100/50 = - 2 Diopter

lens must be diverging

=====

comment before rate in case any doubt, will reply for sure. goodluck